OpenStudy (anonymous):

how do you simplify, (xy^-3)^-2)?

OpenStudy (campbell_st):

will it can be simplified as \[(x)^{-2}\times(y^{-3})^{-2}\] you'll need to use the power of a power law for indices which is \[(x^a)^b = x^{a \times b}\] hope it helps...

OpenStudy (anonymous):

thank you!